where do i start with this equation:
3m-n=-8
m+2n=9


Answer by checkley71(8403)

3M-N=-8 NOW MULTIPLY THE SECOND EQUATION BY -3 & WE GET -3M-6N=-27 NOW ADD
-3M-6N=-27
------------
0M-7N=-35 OR 7N=35 OR N=5 SUBSTITUTING 5 FOR N IN EQUATION - WE GET
3M-5=-8 OR 3M=-3 OR M=-1
PROOF -1+2*5=9 OR -1+10=9 OR 9=9